# A Systematic Survey on Large Language Models for Algorithm Design

### v1

@@ 6.2.6. Other Applications
Other Applications of LLMs extend to a myriad of machine learning tasks, though it is impractical to enumerate them all. Notably, Wong et al. (2024a) capitalize on multi-task learning and the iterative refinement of prompts to foster innovative design approaches. Zeng et al. (2024) integrate LLMs with evolutionary search to develop a heuristic function aimed at efficiently sifting through candidate tensor sets representing the primal tensor network. Furthermore, Guo et al. (2024c) have blazed a trail in employing LLMs to generate novel decision-based adversarial attack algorithms, thus opening up a new diagram for the automatic assessment of model robustness.
@@ 6.3.2. Chemistry, Biology&Physics
In the fields of chemistry and biology, LLMs are predominantly applied in diverse molecule design scenarios. These include general molecule design (Ishida et al., 2024; Bhattacharya et al., 2024), drug molecule design (Ye, 2024), chemical reactions (Ranković and Schwaller, 2023), and catalyst design (Wang et al., 2024a). Physics has been less studied in LLM4AD papers, Pan et al. (2024) demonstrate the capability of LLMs to perform complex analytic calculations in theoretical physics using multi-step prompt templates.
@@ 6.3.3. Mechanics
CodonBERT (Li et al., 2023a) LLMaP mRNA design and optimization
Revisiting-PLMs (Hu et al., 2022) LLMaP protein function prediction
@@ 6.4. Industry
This subsection explores the transformative impact of LLMs on algorithm design across various industries. As we envision the future of manufacturing, LLMs are poised to play a pivotal role. Existing developments, such as those from SymphonyAI  [^1] https://www.symphonyai.com/, focus on enhancing research and investment to refine these models while assessing their performance in terms of safety, bias, and utility. LLMs can be seen as a gateway between humans and machines, facilitating informed decision-making.
For instance, LLMs are instrumental in building a comprehensive intelligent 6G network system (Long et al., 2024), addressing challenges such as low latency, high-frequency bands, high bandwidth, high transmission rates, and overall intelligence (Zhou et al., 2024b). Moreover, LLMs have the potential to revolutionize the telecom industry by streamlining tasks and reducing the need for extensive manpower and engineering expertise. However, it is crucial to address existing challenges to fully realize their potential (Maatouk et al., 2023).
Recently, the application of LLMs in electronic design automation (EDA) has emerged as a promising area of exploration. LLM-based solutions have been developed to enhance interactions with EDA tools (Fang et al., 2024). In particular, within the VLSI design flow, it is essential to verify that the implementation conforms to its specifications. To achieve this, SystemVerilog Assertions (SVA) need to be generated automatically. Generative AI techniques, such as LLMs, can be effectively utilized to facilitate this process (Miftah et al., 2024).
To enhance the reliability and availability of cloud services, conducting root cause analysis (RCA) for cloud incidents is essential. RCACopilot (Chen et al., 2024b) effectively matches incoming incidents with the appropriate incident handlers based on their alert types. It aggregates critical runtime diagnostic information, predicts the incident’s root cause category, and provides an explanatory narrative—all without the need for manual investigations.
@@
Other traditional industrial tasks, such as itinerary planning, can also be enhanced by leveraging the power of LLMs (Xie et al., 2024). LLM-based open-domain urban itinerary planning (Tang et al., 2024b) combines spatial optimization with LLMs to create detailed urban itineraries tailored to users’ requests. This approach extends the scope of itinerary planning by generating personalized urban itineraries based on user requests expressed in natural language.
@@ 7. Challenges of LLMs in Algorithm Design
The application of LLM for algorithm design is not without limitations. This section explores the boundaries of LLMs concerning their scalability, inference costs, and their ability to generate novel ideas specifically in the context of designing algorithms.
@@ 7.1. Scalability
One of the primary limitations of LLMs in algorithm design is their scalability, particularly for handling long inputs and outputs. LLMs are limited by a fixed context window size, restricting the amount of information they can process at once, which is a challenge for complex algorithmic tasks requiring detailed specifications and extensive data. This limitation not only affects their understanding but also the length of the output they can generate, often necessitating multiple iterative queries for comprehensive solutions. Additionally, even when input and output lengths are sufficient, LLMs struggle with comprehension and reasoning over long inputs, impacting their effectiveness in designing complex algorithms. For instance, LLMs tend to solve only low-dimensional problems when used as optimizers (Yang et al., 2023), and they are typically employed to design a function or heuristic component of an algorithm rather than the entire algorithm when used as designers (Liu et al., 2024d).
@@ 7.2. Interpretability
Interpretability is another challenge when using LLMs for algorithm design. It is difficult to trace how specific inputs lead to particular outputs in LLMs. This opacity can be problematic, especially in critical applications, e.g., some industry applications, where understanding the decision-making process is essential for trust and reliability. Efforts to enhance the interpretability of LLMs involve techniques such as feature visualization, model simplification (Liu et al., 2023a), and the study of the contribution of different parts of the data to the model’s behavior (Allen-Zhu and Li, 2023). Despite these efforts, achieving high levels of interpretability without compromising the performance of LLMs remains a key area of ongoing research in the field of artificial intelligence.
7.3. Inference Cost
The deployment of LLMs in algorithm design also faces challenges related to costs, both in terms of training and use. Training domain LLMs requires significant computational resources, often involving extensive GPU or TPU usage over several weeks or even months, which incurs high financial and time costs. Once trained, the inference cost—the computational expense involved in applying the model to new data—also becomes a critical factor, especially when the model is used frequently or requires rapid responses, such as the evolutionary search used in many LLM4AD works (Romera-Paredes et al., 2024). For practical applications in algorithm design, where multiple iterations and tests might be necessary, these costs can accumulate, making the use of LLMs potentially less viable for researchers without substantial computational budgets.
@@ 7.4. Security
7.5. Generation of Novel Ideas
Despite the numerous works on LLM4AD, their ability to produce truly novel ideas is still under scrutiny. LLMs excel at interpolation within the range of their training data but may struggle with extrapolation to new, unseen scenarios. In the context of algorithm design, this means that while LLMs can effectively suggest algorithms or solutions that closely resemble or incrementally improve upon existing ones, they may not be as effective at pioneering radically new algorithmic concepts or methods that deviate significantly from the training data. This limitation is crucial for fields that thrive on innovation and where the generation of novel, out-of-the-box ideas is paramount. The reliance on large volumes of training data also raises concerns about the model’s dependency on historical data, which may bias its outputs towards conventional methods and potentially stifle innovation. While LLM does show the capability of generating new ideas in some tasks (Yao et al., 2024a), how to continuously explore practical novel ideas remains a challenge (Hu et al., 2024).
@@ 8.1. Domain LLM for Algorithm Design
Instead of using a general pre-trained LLM with linguistic and code generation capability, it is worthwhile studying how to train an LLM specifically for automatic algorithm design tasks.
The following aspects can be explored in developing domain LLMs:
  • Training domain LLMs are costly and resource-consuming. On one hand, advanced light fine-tuning techniques, e.g., LORA, could be used for efficient domain adaptation. On the other hand, the size of algorithm LLMs for a specific application can be reduced with the help of domain data and knowledge.
  • Generating and collecting domain data for algorithm design poses challenges. Different from general code generation or linguistic process tasks, there is no large and formatted data specific for algorithm design. Some Github reps on algorithms might be helpful, however, the resources should be scanned and cleaned for algorithm design. Lehman et al. (2024) provide an example of generation domain data for LLMs, where the samples generated during searching are used as the training dataset to fine-tune a pre-trained model resulting in better performance on the target problem.
  • Instead of learning a text and code generation model, how to learn the high-level algorithm design ideas and the algorithmic revision and reasoning ability remain unexplored. The main challenge is how to promote the automatic reasoning capability of LLMs in inference (Zelikman et al., 2022; Wang and Zhou, 2024), which is also recognized as one of the key technologies behind state-of-the-art LLMs, such as OpenAI o1 (OpenAI, 2024).
@@ 8.2. Multi-modal LLM for Algorithm Design
The primary focus of the works discussed in this paper is on utilizing the text comprehension and generation capabilities of LLMs, whether in language, code, or statistics. One distinct advantage of LLMs compared to traditional model-based optimization is their ability to process multi-modal information like humans. Some attempts have been conducted to showcase the advantages of incorporating multi-modal information in algorithm design, as evidenced by Huang et al. (2024c) and Narin (2024). It is anticipated that more principled approaches utilizing multi-modal LLMs will be developed, with further application studies expected in the future.
@@ 8.7. Benchmarking LLM4AD
Benchmarking allows a fair, standardized, and convenient comparison and identifying best practices enabling generating new insights for innovation. While we are glad to witness the emergence of diverse research works and applications, LLM4AD still lacks a benchmarking on either algorithm design test sets for systematic and scientific assessment or a strict pipeline and evaluation criteria on large language models for algorithm design. Several attempts have been made on related topics such as mathematical reasoning (Lin et al., 2024) and planning (Valmeekam et al., 2024). (Veličković et al., 2022) presents an algorithm test set built on some basic algorithms and (McLeish et al., 2024) demonstrates the superior performance of LLM when compared to neural algorithm reasoners. In the future, more benchmarks are expected and they are going to play a crucial role in advancing LLM4AD.
@@ 9. Conclusion
This paper has provided a systematic and up-to-date survey on large language models for algorithm design (LLM4AD), an area that stands at the confluence of artificial intelligence and computational problem-solving. By systematically reviewing a significant corpus of recent literature, this paper not only highlights the current state and evolution of LLM applications in algorithm design but also introduces a novel multi-view taxonomy with comprehensive lists of papers that categorize the roles and functionalities of LLMs in this domain. This taxonomy serves as a framework for both academic and industrial researchers to understand and navigate the complex landscape of algorithm design using LLMs.
The discussion presented in this paper sheds light on the limitations and challenges currently faced by the field, such as issues of scalability, interpretability, and security. Moreover, the proposed future research directions aim to inspire and guide subsequent studies including the need for domain-specific models, multimodal approaches, interactive systems, and comprehensive benchmarking standards.

### v2

@@ 6.2.6. Other Applications
Other Applications of LLMs extend to a myriad of machine learning tasks. Notably, Wong et al. (2024a) capitalize on multi-task learning and the iterative refinement of prompts to foster innovative design approaches. Zeng et al. (2024) integrate LLMs with evolutionary search to develop a heuristic function aimed at efficiently sifting through candidate tensor sets representing the primal tensor network. Furthermore, Guo et al. (2024b) have blazed a trail in employing LLMs to generate novel decision-based adversarial attack algorithms, thus opening up a new diagram for the automatic assessment of model robustness.
@@ 6.3.2. Chemistry, Biology & Physics
In the field of chemistry, LLMs can be applied not only to conventional molecular generation and design (Ishida et al., 2024; Bhattacharya et al., 2024), but also to specialized areas such as drug molecule design (Ye, 2024), chemical reaction prediction and optimization (Ranković and Schwaller, 2023), and catalyst design (Wang et al., 2024a), providing customized solutions. Furthermore, LLMs have also shown promising application prospects in materials discovery, synthesis route planning, green chemistry, and other areas. These studies not only demonstrate the advantages of language models in molecular representation and optimization but also provide new tools and ideas for chemical research.
In biology, LLMs are increasingly being used for tasks such as protein engineering (Lv et al., 2024; Shen et al., 2024), enzyme design (Teukam et al., 2024), and biological sequence analysis (Feng et al., 2024). By combining LLMs with vast amounts of biological data, they can more accurately predict interactions between biological molecules and significantly improve the efficiency of bioinformatics workflows. This has important implications for drug discovery and therapeutic protein design. The unique sequence generation and optimization capabilities of LLMs offer new possibilities for tackling combinatorial optimization challenges in biomacromolecular design.
Although applications in physics are relatively fewer, some emerging work has demonstrated the broad prospects of LLMs. Pan et al. (2024) use multi-step prompt templates to prove that LLMs can perform complex analytical calculations in theoretical physics. Quantum computing algorithm design, physics simulation optimization, and computational methods in condensed matter.
@@ 6.3.3. Mechanics
CodonBERT (Li et al., 2023) LLMaP mRNA Design and Optimization
Revisiting-PLMs (Hu et al., 2022b) LLMaP Protein Function Prediction
@@ 6.4. Industry
This subsection explores the transformative impact of LLMs on algorithm design across various industries. As we envision the future of manufacturing, LLMs are poised to play a pivotal role. Existing developments focus on enhancing research and investment to refine these models while assessing their performance in terms of safety, bias, and utility. LLMs can be seen as a gateway between humans and machines, facilitating informed decision-making.
For instance, LLMs are instrumental in building a comprehensive intelligent 6G network system (Long et al., 2024), addressing challenges such as low latency, high-frequency bands, high bandwidth, high transmission rates, and overall intelligence (Zhou et al., 2024b). Moreover, LLMs have the potential to revolutionize the telecom industry by streamlining tasks and reducing the need for extensive manpower and engineering expertise. However, it is crucial to address existing challenges to fully realize their potential (Maatouk et al., 2024).
Recently, the application of LLMs in Electronic Design Automation (EDA) has emerged as a promising area of exploration. LLM-based solutions have been developed to enhance interactions with EDA tools (Fang et al., 2024). In particular, within the VLSI design flow, it is essential to verify that the implementation conforms to its specifications. To achieve this, system verilog assertions need to be generated automatically. Generative AI techniques, such as LLMs, can be effectively utilized to facilitate this process (Miftah et al., 2024).
To enhance the reliability and availability of cloud services, conducting root cause analysis (RCA) for cloud incidents is essential. RCACopilot (Chen et al., 2024b) effectively matches incoming incidents with the appropriate incident handlers based on their alert types. It aggregates critical runtime diagnostic information, predicts the incident’s root cause category, and provides explanatory results without the need for manual investigations.
@@
Other traditional industrial tasks, such as itinerary planning, can also be enhanced by leveraging the power of LLMs (Xie et al., 2024). LLM-based open-domain urban itinerary planning (Tang et al., 2024b) combines spatial optimization with LLMs to create detailed urban itineraries tailored to users’ requests. This approach extends the scope of itinerary planning by generating personalized urban itineraries based on user requests in natural language.
@@ 7.1. Scalability
One of the primary limitations of LLMs in algorithm design is their scalability, particularly for handling long inputs and outputs. LLMs are limited by a fixed context size, restricting the amount of information they can process at once, which is a challenge for complex algorithmic tasks requiring detailed specifications and extensive content. Additionally, even when input and output lengths are sufficient, LLMs struggle with comprehension and reasoning over long inputs, impacting their effectiveness in designing complex algorithms. For instance, LLMs tend to solve only low-dimensional problems when used as optimizers (Yang et al., 2024), and they are employed to design a function or heuristic component of an algorithm rather than the entire algorithm when used as designers (Liu et al., 2024d).
@@ 7.2. Interpretability
Interpretability is another challenge when using pre-trained LLMs for algorithm design. It is difficult to trace how specific inputs lead to particular outputs in LLMs due to their black-box nature. This opacity can be problematic, especially in critical applications, e.g., some industry applications, where understanding the decision-making process is essential for trust and reliability. Efforts to enhance the interpretability involve techniques such as feature visualization, model simplification (Liu et al., 2023a), and the study of the contribution of different parts of the data to the model’s behavior (Allen-Zhu and Li, 2023). Despite these efforts, achieving principled interpretability without compromising the performance of LLMs remains a key area of ongoing research.
@@ 7.3. Security
7.4. Cost
The utilization of LLMs in algorithm design also faces challenges related to costs, both in terms of model training and usage. Training domain LLMs requires significant computational resources, often involving extensive GPU or TPU usage over several weeks or even months, which incurs high financial and time costs. Once trained, the inference cost—the computational expense involved in applying the model to algorithm design—also becomes a critical factor, especially when the model is used frequently or requires rapid responses, such as the evolutionary search used in many LLM4AD works (Romera-Paredes et al., 2024). For practical applications in algorithm design, where multiple iterations and tests might be necessary, these costs can accumulate, making the use of LLMs potentially less viable for researchers without substantial computational budgets.
7.5. Idea Innovation
Despite the numerous works on LLM4AD, their ability to produce truly novel ideas is still under scrutiny. LLMs excel at interpolation within the range of their training data but may struggle with extrapolation to new, unseen scenarios. In the context of algorithm design, this means that while LLMs can effectively suggest algorithms or solutions that closely resemble or incrementally improve upon existing ones, they may not be as effective at pioneering radically new algorithmic concepts or methods. The reliance on large volumes of training data also raises concerns about the model’s dependency on historical data, which may bias its outputs towards conventional methods and potentially stifle innovation. While LLM does show the capability of generating new ideas in some tasks (Yao et al., 2024a), how to continuously explore practical novel ideas remains a challenge (Hu et al., 2024).
@@ 8.1. Domain LLM for Algorithm Design
Instead of using a general pre-trained general-purpose LLM, it is worthwhile studying how to train an LLM specifically for automatic algorithm design tasks. The following aspects can be explored in developing domain LLMs: 1) Training domain LLMs are costly and resource-consuming. On the one hand, advanced light fine-tuning techniques (e.g., LoRA (Hu et al., 2022a)) could be used for efficient domain adaptation. On the other hand, the size of algorithm LLMs for a specific application can be reduced with the help of domain data and knowledge. 2) Generating and collecting domain data for algorithm design poses challenges. Unlike general code generation or linguistic process tasks, there is no large and formatted data specific for algorithm design. Some Github repositories on algorithms might be helpful, but the resources should be scanned and cleaned for algorithm design. Lehman et al. (2024) provide an example of generation domain data for LLMs, where the samples generated during searching are used as the training dataset to fine-tune a pre-trained model resulting in better performance on the target problem. 3) Instead of learning a text and code generation model, how to learn the algorithm development ideas and the algorithmic reasoning ability remain unexplored (Wang and Zhou, 2024; OpenAI, 2024).
@@ 8.2. Multi-modal LLM for Algorithm Design
The primary focus of existing LLM4AD works is on utilizing the text comprehension and generation capabilities of LLMs, whether in language, code, or statistics. One advantage of LLMs compared to traditional model-based optimization is their ability to process multi-modal information like humans, which is rarely studied. Some attempts have been conducted to showcase the advantages of incorporating multi-modal information in algorithm design, as evidenced by Huang et al. (2024c) and Narin (2024). It is anticipated that more methods and applications utilizing multi-modal LLMs will be developed in the future.
@@ 8.7. Benchmarking LLM4AD
Benchmarking allows for a fair, standardized, and convenient comparison, which can identify best practices and enable generating new insights for innovation. While we are glad to witness the emergence of diverse research works and applications, LLM4AD still lacks a benchmarking on either algorithm design test sets for systematic and scientific assessment or a strict pipeline and evaluation criteria on large language models for algorithm design. Several attempts have been made on related topics such as mathematical reasoning (Lin et al., 2024) and planning (Valmeekam et al., 2024). (Veličković et al., 2022) presents an algorithm test set built on some basic algorithms and (McLeish et al., 2024) demonstrates the superior performance of LLM when compared to neural algorithm reasoners. In the future, more benchmarks are expected and they are going to play a crucial role in advancing LLM4AD.
@@ 9. Conclusion
This paper has provided a systematic and up-to-date survey on large language models for algorithm design (LLM4AD). By systematically reviewing a significant corpus of main contributions in this emerging research field, this paper not only highlights the current state and evolution of LLM applications in algorithm design but also introduces a novel multi-dimensional taxonomy with comprehensive lists of papers that categorize the roles and functionalities of LLMs in this domain. This taxonomy serves as a framework for both academic and industrial researchers to understand and navigate the complex landscape of algorithm design using LLMs. We have also discussed the limitations and challenges currently faced by the field, such as issues of scalability, interpretability, and security. Moreover, the proposed future research directions aim to inspire and guide subsequent studies.